OpenStudy (anonymous):

HELP PLEASE!!! Write the function that models the relationship. Find z when x = 4 and y = 8. z varies jointly with x and y. When x = 2 and y = 2, z = 7. I'm a bit behind in alg 2 and I have to complete a study guide for a test tomorrow and I'm stuck on this question! Help would be much appreciated! (I know that the answer is z = 7/4xy, but I don't know how to get there.)

OpenStudy (***[isuru]***):

do u know what is mean by joint variation ?

OpenStudy (anonymous):

it's when one quantity varies directly with 2 or more quantities?

OpenStudy (***[isuru]***):

yeah... then \[z \ \ \ \alpha \ \ xy\]

OpenStudy (anonymous):

since you have the equation jus fill in what you have

OpenStudy (***[isuru]***):

so.. u can write it as z = k xy

OpenStudy (***[isuru]***):

u've know the values of x and y when z = 7

OpenStudy (***[isuru]***):

z = 7 x = 2 y = 2

OpenStudy (anonymous):

so it would be 7 = k(2)(2)?

OpenStudy (anonymous):

well, k(4)

OpenStudy (***[isuru]***):

and u know that z = k xy use the above values and find k

OpenStudy (***[isuru]***):

yes!

OpenStudy (***[isuru]***):

then k will equal to.. 7 = 4k k = 7/4

OpenStudy (***[isuru]***):

now... write the equation z = k xy substituting the value for k and that;s the equation u want...!!

OpenStudy (anonymous):

wow okay that was so much easier than i thought! thank you so much! ^_^

OpenStudy (***[isuru]***):

now.. to find z when x= 4 and y = 8 plug those value to the equation u got.. it;s that simple :)

OpenStudy (***[isuru]***):

u r welcome!